# A modular perspective to the jet suppression from a small to large radius in very high transverse momentum jets

### v3

@@ II.1 Exploring the MARTINI and AdS/CFT models
Simulations of Pb-Pb collisions are carried out using the MATTER module for both vacuum and medium showers, integrated with three distinct time-ordered parton shower models: LBT, MARTINI and AdS/CFT. The transition between different energy loss modules is performed independently for each parton, with the switching parameter $Q_{0}$ set at 2 GeV. Both the MARTINI [32] and the AdS/CFT [33] modules are developed to handle the low virtuality phase and simulate the energy loss once the parton is transferred from the MATTER. We replace the successful LBT module with the MARTINI, to carry out the simulations for the same conditions [49] and compare the jet-${R_{AA}}$ with the experimental data from the CMS [35]. Similarly, the calculations are done by replacing the LBT with AdS/CFT module and compared with the experimental data.
This section presents the results for low virtuality parton evolution models, where the inclusive jet-${R_{AA}}$ is calculated for three setups, MATTER coupled with LBT, MATTER coupled with MARTINI, and MATTER coupled with AdS/CFT. The simulations are altogether staged in contrast with the experimental data from the CMS [35], for Pb-Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02}$ TeV and for a range of jet cone sizes up to an order of 1 and $p_{T}^{jet}$ range up to 1 TeV. The measurements for different models are now manifest in Fig. 4, which shows the jet-${R_{AA}}$ for ${|y_{jet}|<2}$ and jet cone radius $R$ = 0.2, $R$ = 0.3, $R$ = 0.4, $R$ = 0.6, $R$ = 0.8, and $R$ = 1.0 over a range of jet-${p_{T}}$ from 300 GeV up to 1 TeV in comparison with the experimental data from the CMS [35]. The predictions made by the JETSCAPE are consistent even as we advance to the larger area jet cones. The JETSCAPE predicts marginally more suppression from a small to intermediate to large radius for all the modular combinations discussed above. The results predicted by the (MATTER + LBT) and (MATTER + MARTINI) are congruous with the experimental observations throughout the ${p_{T}}$ range and across the different jet radii. While the (MATTER + AdS/CFT) model shows more suppression at high ${p_{T}}$ region as compared to the other modular combinations. At very high transverse momentum, the virtuality of the dominant partons is very high and the coherence effect comes into play. The coherence effect significantly reduces the interaction strength of the parton with the medium. However, AdS/CFT model is known for its strong coupling framework, which results in more significant energy loss for partons traversing the QGP medium. This leads to stronger suppression of jet-${R_{AA}}$ compared to the perturbative approaches used by LBT and MARTINI at high ${p_{T}}$ range.
@@
Figure 5: The double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet radius for inclusive jets for (${\mathrm{300~{}GeV}\leq p_{T,jet}\leq\mathrm{400~{}GeV}}$) (left panel), (${\mathrm{400~{}GeV}\leq p_{T,jet}\leq\mathrm{500~{}GeV}}$) (centre panel), and (${\mathrm{500~{}GeV}\leq p_{T,jet}\leq\mathrm{1~{}TeV}}$) (right panel), in the most central(0-10%) Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02\rm~{}TeV}$ for different jet radii with ${|y_{jet}|<2}$ and a minimum track requirement of ${p^{\mathrm{track}}_{\mathrm{T}}>}$ 0.5 GeV. The plot shows the comparison between the models by magenta, red and blue markers for (MATTER+LBT), (MATTER+MARTINI), and (MATTER+AdS/CFT), respectively. The predictions are in comparison with CMS data [35] shown with the black marker.
@@ II.2 Jet radius (${R}$) and jet-${p_{\mathrm{T}}}$ dependence of the ${R_{AA}}$
In this section, we emphasize the significant contribution from the hydrodynamic medium response already highlighted in previous studies [8, 56], which is observed in the jet-${R_{\mathrm{AA}}}$ as a function of jet radius. Through radius-dependent studies, we develop a clear picture of the role played by radiation and collisions in energy loss.
A recent study by the CMS collaboration [35] enlightens the rigorous comparisons of predictions from quenched jet event generators like HYDJET++ [57], PYQUEN [58] and theoretical models used to replicate relativistic heavy ion collisions to the experimental data for the jet-${R_{\mathrm{AA}}}$. The article concludes with the impression that although most state-of-the-art models have progressed, significant uncertainty remains for the large-area jets.
@@
This quest is executed by calculating the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet radius which is in comparison with the CMS data for the most central (0-10%) Pb-Pb collisions over a range of jet-${p_{T}}$ from 300 GeV up to 1 TeV. The results are further sub-categorized by three jet-${p_{T}}$ intervals as shown in Fig. 5.
@@
For ${\mathrm{300~{}GeV}\leq p_{T,jet}\leq\mathrm{400~{}GeV}}$ interval, Fig. 5 (left panel) shows the predictions by the different combination of energy loss models. However, we note that there is only one measurement done by the CMS [35] in this ${p_{T}}$ interval, i.e., for $R$ = 0.3. We observe a consistent and nearly identical trend across all three models, i.e.,(MATTER + LBT), (MATTER + AdS/CFT), and (MATTER + MARTINI), each predicting a slight enhancement in the double ratio and showing good agreement with the data. This also indicates that the energy loss of the partons in the intermediate ${p_{T}}$ range throughout the various jet cone sizes is comparable among LBT, MARTINI, and AdS/CFT.
For ${\mathrm{400~{}GeV}\leq p_{T,jet}\leq\mathrm{500~{}GeV}}$ interval, in Fig. 5 (center panel), we observe that all models predict a similar trend, showing a consistent, slight increase that aligns well with the observed data. The (MATTER + AdS/CFT) model, in particular, predicts the jet-${R_{\mathrm{AA}}}$ within approximately ${15\%}$ of the data, providing a closer match than the other models. In the extreme high $p_{T,jet}$ region, that is for ${\mathrm{500~{}GeV}\leq p_{T,jet}\leq\mathrm{1~{}TeV}}$, in Fig. 5 (right panel), shows that all the models significantly over-predict the jet-${R_{\mathrm{AA}}}$ double ratio, with the crème de la crème description provided by the (MATTER + AdS/CFT) model up to ${10\%}$, as it also harmonises with the data within the uncertainties. However, (MATTER + MARTINI) slightly overpredicts though it follows the same general pattern as for the other $p_{T,jet}$ regions. All the models predict that the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) should increase with R, which leads to the same pattern throught all the $p_{T,jet}$ regions. This alignment with data strengthens confidence in the theoretical descriptions of jet quenching mechanisms and the influence of jet radius, supporting the models’ assumptions about energy loss and medium effects on jets.
Altogether, the JETSCAPE framework convincingly describes the full evolution of the parton shower by adopting a virtuality-based multi-stage approach for energy loss. We have also demonstrated that MARTINI when combined with MATTER, reveals the effect of using the (2+1)D MUSIC hydrodynamic model for the soft background on the results of the hard observable, the jet-${R_{\mathrm{AA}}}$.
@@
We also measure the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet-${p_{T}}$ for (MATTER + LBT) model in Fig. 6. We observe a non-monotonous rise in the double ratio for jet-${p_{T}}\leq$ 200 GeV, followed by a fall in the intermediate ${p_{T}}$ to high ${p_{T}}$ region. The double ratio of larger radius jets is higher as compared to the smaller radius jets. As the jet is scattered in the medium, the final state partons of the medium interacting with the jet are also considered a constituent of the jet. As more and more gluons encapsulated into the larger jet cone, the energy lost by the jets is partially gained as the area of the jet cone increases.
@@
Figure 6: The double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet-${p_{T}}$ for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02\rm~{}TeV}$ for different jet radii.
@@ III Conclusion
In this paper, we present the jet-${R_{AA}}$ predictions from the JETSCAPE framework incorporating the (2+1)D MUSIC model for viscous hydrodynamic evolution and compare with the ATLAS data for Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02}$ TeV in jet transverse momentum interval ${\mathrm{100~{}GeV}\leq p_{T,jet}\leq\mathrm{1~{}TeV}}$ for jets of radius R = 0.4. These results show that the (2+1)D MUSIC model along with (MATTER + LBT) is successful in describing the experimental data even at higher jet-${p_{T}}$ for the most central (0-10%) collisions. This work also elucidates the predictions made by low virtuality-based evolution models like MARTINI and AdS/CFT in a hydrodynamic medium generated by the (2+1)D MUSIC when compared to the CMS data. We observe more suppression in jet-${R_{AA}}$ in the low to intermediate ${p_{T}}$ region, providing a better description of the observed data.
We advance the current JETSCAPE calculations and compare with the data of wider jet cones ranging from $R$ = 0.2 to $R$ = 1.0 measured by the CMS detector for Pb+Pb collisions at ${\sqrt{s_{\mathrm{NN}}}=5.02}$ TeV in jet transverse momentum interval ${\mathrm{300~{}GeV}\leq p_{T,jet}\leq\mathrm{1~{}TeV}}$. It suggests that over time, jets are increasingly able to recover more energy within larger radii. This can indicate that as jets propagate through the medium, more medium-induced radiation or scattered particles are incorporated within larger-radius jets.
In this study, we have also shown the dependence of jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) on the jet-${p_{T}}$, which provides critical insights into the behavior of jet suppression in higher $p_{T}^{jet}$ region. The analysis reveals consistent trend for (MATTER + LBT), which shows that the double ratio is higher for jets with larger radii than for those with smaller radii. This suggests that jets with smaller radii experience a greater relative suppression effect in the high-${p_{T}}$ region, while larger-radius jets recover more energy within the broader cone, leading to reduced suppression.
The predictions made by the JETSCAPE framework indicate that the jets with larger radii and high ${p_{T}}$, retain a greater fraction of the initial hard parton momentum. As the jet cone expands, it captures more of the radiated energy and medium-scattered partons, which would otherwise be lost in smaller jets. This broader structure allows larger-radius jets to partially recover the energy lost through quenching interactions within the medium, resulting in a higher ${R_{\mathrm{AA}}}$ and reduced suppression compared to smaller-radius jets.
This work highlights the current standing of the JETSCAPE framework to describe the energy loss and medium response phenomenon for broad-area jet cones and for very high transverse momentum jets.
@@
We express our gratitude to the JETSCAPE Collaboration for making the state-of-the-art framework publicly available for extensive research use. The authors would like to especially acknowledge the members of the JETSCAPE Collaboration, Yasuki Tachibana, Abhijit Majumder, Amit Kumar, and Chun Shen for their useful discussion and valuable feedback. The authors would also like to thank IIT Mandi Param-Himalaya computing facility, Goa University Param computing facility, SPAS local cluster facility, and seed money grant support.

### v4

@@ II.1 Exploring the MARTINI and AdS/CFT models
Simulations of Pb-Pb collisions are carried out using the MATTER model for both vacuum and medium showers, integrated with three distinct time-ordered parton shower models: LBT, MARTINI and AdS/CFT. The transition between different energy loss models is performed independently for each parton, with the switching parameter $Q_{\mathrm{SW}}$ set at 2 GeV. Both the MARTINI [32] and the AdS/CFT [33] models are developed to handle the low virtuality phase and simulate the energy loss once the parton is transferred from the MATTER. We replace the successful LBT model with the MARTINI, to carry out the simulations for the same conditions [49] and compare the jet-${R_{AA}}$ with the experimental data from the CMS [35]. Similarly, the calculations are done by replacing the LBT with AdS/CFT model and compared with the experimental data.
This section presents the results for low virtuality parton evolution models, where the inclusive jet-${R_{AA}}$ is calculated for three setups, MATTER coupled with LBT, MATTER coupled with MARTINI, and MATTER coupled with AdS/CFT. The simulations are altogether staged in contrast with the experimental data from the CMS [35], for Pb-Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV and for a range of jet cone sizes up to an order of 1 and $p_{T}^{jet}$ range up to 1 TeV. The measurements for different models are now manifest in Fig. 4, which shows the jet-${R_{AA}}$ for ${|y_{jet}|<2}$ and jet cone radius $R$ = 0.2, $R$ = 0.3, $R$ = 0.4, $R$ = 0.6, $R$ = 0.8, and $R$ = 1.0 over a range of jet-${p_{T}}$ from 300 GeV up to 1 TeV in comparison with the experimental data from the CMS [35]. The predictions made by the JETSCAPE are consistent even as we advance to the larger area jet cones. The JETSCAPE predicts marginally more suppression from a small to intermediate to large radius for all the modular combinations discussed above. The results predicted by the (MATTER + LBT) and (MATTER + MARTINI) are congruous with the experimental observations throughout the ${p_{T}}$ range and across the different jet radii. While the (MATTER + AdS/CFT) model shows more suppression at high ${p_{T}}$ region as compared to the other modular combinations. At very high transverse momentum, the virtuality of the dominant partons is very high and the coherence effect comes into play. The coherence effect significantly reduces the interaction strength of the parton with the medium. However, AdS/CFT model is known for its strong coupling framework, which results in more significant energy loss for partons traversing the QGP medium. This leads to stronger suppression of jet-${R_{AA}}$ compared to the perturbative approaches used by LBT and MARTINI at high ${p_{T}}$ range.
@@
Figure 5: The double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet radius for inclusive jets for (${\mathrm{300~{}GeV}\leq p_{T}^{jet}\leq\mathrm{400~{}GeV}}$) (left panel), (${\mathrm{400~{}GeV}\leq p_{T}^{jet}\leq\mathrm{500~{}GeV}}$) (centre panel), and (${\mathrm{500~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$) (right panel), in the most central(0-10%) Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for different jet radii with ${|y_{jet}|<2}$. The plot shows the comparison between the models by magenta, red and blue markers for (MATTER+LBT), (MATTER+MARTINI), and (MATTER+AdS/CFT), respectively. The predictions are in comparison with CMS data [35] shown with the black marker.
@@ II.2 Jet radius (${R}$) and jet-${p_{\mathrm{T}}}$ dependence of the ${R_{AA}}$
In this section, we emphasize the significant contribution from the hydrodynamic medium response, as already highlighted in previous studies [8, 55], which is observed in the jet-${R_{\mathrm{AA}}}$ as a function of jet radius. Through radius-dependent studies, we develop a clear picture of the role played by radiation and collisions in energy loss.
A study by the CMS collaboration [35] enlightens the rigorous comparisons of predictions from quenched jet event generators like HYDJET++ [56], PYQUEN [57] and theoretical models used to replicate relativistic heavy-ion collisions to the experimental data for the jet-${R_{\mathrm{AA}}}$. The article concludes with the impression that although most state-of-the-art models have progressed, there is still considerable uncertainty regarding large-area jets.
@@
This quest is executed by calculating the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of jet radius which is in comparison with the CMS data for the most central (0-10%) Pb-Pb collisions. The double ratio acts as a key indicator for assessing the dependence of jet energy loss on jet-$R$. When the ratio falls below unity, it indicates greater suppression of jets with larger R, while close to unity suggests no R-dependence or a balance of competing effects. Conversely, a ratio exceeding unity implies reduced suppression for jets with larger R, as demonstrated in the previous section. The results are obtained over a range of jet-${p_{T}}$ from 300 GeV up to 1 TeV and are further sub-categorized into three distinct jet-${p_{T}}$ intervals as shown in Fig. 5.
@@
For ${\mathrm{300~{}GeV}\leq p_{T}^{jet}\leq\mathrm{400~{}GeV}}$ interval, Fig. 5 (left panel) shows the predictions by the different combination of energy loss models. However, we note that there is only one measurement done by the CMS [35] in this ${p_{T}}$ interval, i.e., for $R$ = 0.3. We observe a consistent and nearly identical trend across all three models, i.e., (MATTER + LBT), (MATTER + AdS/CFT), and (MATTER + MARTINI), each predicting a minimal enhancement in the double ratio and showing good agreement with the data. This also indicates that the energy loss of the partons throughout the various jet cone sizes is comparable among LBT, MARTINI, and AdS/CFT.
For ${\mathrm{400~{}GeV}\leq p_{T}^{jet}\leq\mathrm{500~{}GeV}}$ interval, in Fig. 5 (center panel), we observe that the models (MATTER + LBT) and (MATTER + AdS/CFT) predict a similar trend, showing a consistent, marginal increase that aligns well with the observed data. The (MATTER + AdS/CFT) model, in particular, predicts the jet-${R_{\mathrm{AA}}}$ within ${15\%}$ of the experimental data, providing a closer match than the other models. However, the (MATTER + MARTINI) model gives double ratio values near unity in the intermediate ${p_{T}}$ range, showing a negligible R-dependence, which stands in contrast to the CMS observation.
In the extreme high $p_{T}^{jet}$ region, specifically for ${\mathrm{500~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$, shown in Fig. 5 (right panel), the experimental data reveal that double ratio is less than unity for intermediate R. This may indicate that highly boosted jets, which possess more intricate jet substructures, undergo marginally more suppression. However, all the models moderately overestimate the jet-${R_{\mathrm{AA}}}$ double ratio in this high $p_{T}^{jet}$ region. Among them, the (MATTER + AdS/CFT) model provides the closest agreement, deviating by only up to ${10\%}$, and also harmonises with the data within the uncertainties. In contrast, (MATTER + MARTINI) over-predicts the ratio although it follows the same general pattern observed in other $p_{T}^{jet}$ regions.
All the models predict that the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) increases with the jet-$R$, showing a systematic pattern across all the $p_{T}^{jet}$ regions. This agreement with data boosts confidence in the theoretical descriptions of jet quenching and the impact of jet radius on energy loss and medium effects.
Altogether, the JETSCAPE framework convincingly describes the complete evolution of the parton shower by adopting a virtuality-based multi-stage approach for energy loss. Across the entire $p_{T}^{jet}$ regions, the radius-dependent double ratio predictions from the LBT, MARTINI and AdS/CFT models exhibit distinct jet energy loss mechanisms. MARTINI, which combines both radiative and collisional energy loss processes, shows moderate dependence on jet-$R$, with a gradual rise in double ratio as larger radii recover more radiated energy. However, the model demonstrates greater sensitivity at higher $p_{T}$ region, where the increased contribution from gluon radiation leads to more significant changes in the double ratio. LBT exhibits a more consistent dependence on radius as compared to other two models, with its inclusion of medium recoil and dynamic effects from the QGP [58]. This allows LBT to more effectively recover the energy lost outside the smaller jet cones as the jet-$R$ increases. However, AdS/CFT, based on a strongly coupled medium, displays minimal radius dependence, as energy is predominantly dissipated into the medium with limited recovery within the jet cone. This difference highlights the varying assumptions regarding jet-medium interactions through all $p_{T}^{jet}$ regions, with LBT offering a more suitable framework for describing energy loss recovery.
@@
We extend our analysis by calculating the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of $p_{T}^{jet}$ for (MATTER + LBT) model in Fig. 6. We observe a non-monotonous rise in the double ratio for $p_{T}^{jet}\leq$ 200 GeV, followed by a fall in the intermediate ${p_{T}}$ to the high ${p_{T}}$ region. The double ratio of larger radius jets is higher as compared to the smaller radius jets [59, 60]. As the jet is scattered in the medium, the final state partons from the medium that interact with the jet are also considered as the constituents of the jet, with many being emitted at large angles relative to the jet axis. As a result, increasing the jet radius enables the recovery of energy that would otherwise be lost outside the small jet-cone. Furthermore, the medium’s response to the jet, often referred to as the wake effect [61], contributes by redirecting particles back into the jet cone. This process can lead to an increment of ${R_{\mathrm{AA}}}$ for larger R as compared with smaller R-cones. Therefore, as more gluons are captured within the larger jet cone, the energy lost is partially gained as the area of the jet cone increases.
@@
Figure 6: The double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of $p_{T}^{jet}$ for inclusive jets in the most central (0-10%) Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for different jet radii.
@@ III Conclusion
In this paper, we present the jet-${R_{AA}}$ predictions from the JETSCAPE framework, which incorporates the (2+1)D MUSIC model for viscous hydrodynamic evolution and compare them with the ATLAS data for Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV in jet transverse momentum interval ${\mathrm{100~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$ for jets of radius R = 0.4. These results show that the (2+1)D MUSIC model along with (MATTER + LBT) is successful in describing the experimental data even at higher jet-${p_{T}}$ for the most central (0-10%) collisions. This work also elucidates the predictions made by low virtuality-based evolution models like MARTINI and AdS/CFT in a hydrodynamic medium generated by the (2+1)D MUSIC, when compared to the CMS data. We observe marginally more suppression in jet-${R_{AA}}$ in the low to intermediate ${p_{T}}$ region.
We advance the current JETSCAPE calculations and compare with the data of wider jet cones ranging from $R$ = 0.2 to $R$ = 1.0 measured by the CMS detector for Pb+Pb collisions at ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV in jet transverse momentum interval ${\mathrm{300~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$. It suggests that, jets are increasingly able to recover more energy within larger radii. This can indicate that as jets propagate through the medium, more medium-induced radiation or scattered particles are incorporated within larger-radius jets.
In this study, we have also shown the dependence of jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) on the jet-${p_{T}}$. The analysis provides a consistent trend for (MATTER + LBT), which shows that the double ratio is higher for jets with larger radii than for those with smaller radii. This suggests that, the larger-radius jets retain a greater fraction of the initial hard parton momentum by capturing more radiated energy within the broader cone. This leads to the reduced suppression and a higher ${R_{\mathrm{AA}}}$ compared to smaller-radius jets.
This work highlights the potential of the JETSCAPE framework to describe the energy loss and medium response phenomenon for broad-area jet cones and for very high transverse momentum jets.
@@
We express our gratitude to the JETSCAPE Collaboration for making the state-of-the-art framework publicly available for extensive research use. The authors would like to specifically thank the members of the JETSCAPE Collaboration, Yasuki Tachibana, Abhijit Majumder, Amit Kumar, and Chun Shen for their insightful discussion and valuable feedback. The authors would also like to thank IIT Mandi Param Himalaya computing facility under National Supercomputing Mission, DST, SPS local cluster facility, and the IIT Mandi SRIC seed grant support (Ref. No. IITM/SG/2024/01-2348).